How many ways can a president and a vice president be selected in a class of 25 students
Found 2 solutions by jim_thompson5910, ewatrrr:
Answer by jim_thompson5910(35256)

There are 25 choices for the president and 25-1 = 24 choices for the VP. So there are 25*24 = 600 different ways.

Answer by ewatrrr(24785)


 
Hi,
How many ways can a president and a vice president be selected in a class of 25 students
Order important, Permutation used
25P2 = 25!/23! = 25·24 = 600 ways